Date: Sun, 2 May 2004 05:14:32 -0700 (MST)
From: Ettore Pedretti <>
Subject: End of run report 

Hello, All

IOTA has been behaving very well... I think I solved the problem causing
the crash of DAQ and necessary reboot of the VME rack. It was caused by
the state machine staying in quadrant mode while loading the data readout
circuit. I now reset the state machine to "IDLE" state before leaving the
quadrant program. Never had a single crash since... The modified
software is separated from the official software release and can be run by
invoking the script observe_ep_g3.2.

I found the raster scan, written by John for the fiber explorer, very
useful and time saving.

Now a list of problems:

Lots of laser light in one of the beams can sometimes confuse the star
tracker. Has a baffle moved?

The movable LED targets installed by Marc in front of the exit mirrors
have moved, so, do not use them until Marc fixes the problem.

SD1 Yaw is totally stuck. Fortunately LD1 mirror is in a good position
just now. This problem must be addressed soon since it is potentially

Sd1 acts erratically. It sometimes does not find home switch. The problem
seems random but it is getting worse.

South piezo motor: one of the axis works intermittently during alignment.

[About the recent hack]